OpenStudy (anonymous):

A baseball diamond's home plate has three right angles. The other two are congruent. Find their measure.

OpenStudy (tkhunny):

Exterior angles ALWAYS add to 360º.

OpenStudy (anonymous):

Well, first off. This is the homeplate looks like.  The three right angles are at the bottom right, the bottom left, and the top. As for the equation for the sum of the angles in a polygon, where this one is a pentagon, 180(n-2) where n is the number of sides. 180(5-2) 180(3) =540 So, we know that 3 of the angles are 90 degrees, and the other two are equal to each other. 2x+3(90)=540 2x+270=540 -270 from both sides 2x=270 divide by 2 x= 135 degrees.